OpenStudy (anonymous):

During the first part of a trip, a canoeist travels 89 miles at a certain speed. The canoeist travels 23 miles on the second part of the trip at a speed 5 mph slower. The total time for the trip is 3 hrs. What was the speed on each part of the trip?

OpenStudy (anonymous):

\[\frac{89}{a}+\frac{23}{a-5}=3\]

OpenStudy (anonymous):

Solve for a

OpenStudy (anonymous):

a=14

OpenStudy (anonymous):

damn it is 4 hours

OpenStudy (anonymous):

there should be another answer close to 38

OpenStudy (anonymous):

36.25 ??

OpenStudy (anonymous):

89(a−5)+23(a)=4(a)(a−5)

OpenStudy (anonymous):

I still can not get the answer.

OpenStudy (anonymous):

\[\approx29.19\] hardcore quadratics :)

OpenStudy (anonymous):

that is for the first part what is the second part

OpenStudy (anonymous):

let's see what computer tell us

OpenStudy (anonymous):

3.81

OpenStudy (anonymous):

3.85 and 38.48 ??

OpenStudy (anonymous):

yes, but you only take 38 because other one does not make sense.

OpenStudy (anonymous):

the first answer was right 29.19 so the second part is 38.48

OpenStudy (anonymous):

you also have to say what (a-5) is

OpenStudy (anonymous):

well that is not right

OpenStudy (anonymous):

first was right second answer was wrong

OpenStudy (anonymous):

a=38.48 a-5=33.48 a=first part b= second part

OpenStudy (anonymous):

B was 24.19